What is the accumulated sum of each of the following streams of
payments?
a. $500 a year for 10 years compounded annually at 5 percent

Answer by stanbon(75887)

What is the accumulated sum of each of the following streams of
payments?
a. $500 a year for 10 years compounded annually at 5 percent
-----------
A(t) = P(1+(r/n))^(nt)
----------------------------
A(10) = 500(1+(0.05/1)^(1*10)
A(10) = 500(1.05)^10
A(10) = 500*1.6289
A(10) = $814.45
====================
Cheers,
Stan H.